# Real-Time Coordination of Integrated Transmission and Distribution Systems: Flexibility Modeling and Distributed NMPC Scheduling

### 01 Feb 2024 v1

@@ Abstract
This paper proposes a real-time distributed operational architecture to efficiently coordinate integrated transmission and distribution systems (ITD). At the distribution system level, the distribution system operator (DSO)computes the aggregated flexibility of all controllable devices by power-energy envelopes and provides them to the transmission system operator (TSO). At the transmission system level, a distributed nonlinear model predictive control (nmpc)approach is proposed to coordinate the economic dispatch of multiple TSOs, considering the aggregated flexibility of all distribution systems. The subproblems of the proposed approach are associated with different TSOs and individual time periods. In addition, the aggregated flexibility of controllable devices in distribution networks is encapsulated, re-calculated, and communicated through the power-energy envelopes, significantly reducing computational costs and eliminating redundant information exchanges between TSOs and DSOs for privacy and security preservation. The extensive numerical experiment demonstrates the scalability and the computational efficiency of the proposed distributed operational architecture.
@@
This work was supported in part by the BMBF-project ENSURE II with grant number 03SFK1F0-2, in part by the Swiss National Science Foundation (SNSF) under the NCCR Automation project, grant agreement 51NF40_180545, and in part by the Swiss Federal Office of Energy’s “SWEET” programme and performed in the PATHFNDR consortium.
@@ I Introduction
Multiperiod dispatch problems for ITD systems usually couple individual steady-state optimal power flow (OPF)optimization problems over multiple time periods [4, 5, 6]. The coupling constraints include the generator ramping limits, the model of distributed energy storage systems (ESS), and other time-dependent constraints to consider the controllable devices with time-variant properties. However, it is still a challenge to solve a multiperiod AC optimal power flow (MPOPF)for ITD systems. On the one hand, the AC OPF is generally NP-hard [7], and the complexity of solving an MPOPF is further magnified by the intercoupling of subsequent time periods [5]. On the other hand, collecting necessary and realistic data from multiple stakeholders (i.e., TSOs and DSOs), including grid topology, load profiles, and other sensitive information regarding consumer behaviors, is either not preferred or restricted by regulations [3]. To address these challenges and achieve an efficient operation of the overall ITD, recent research analyzed the determination of the aggregated flexibility of the controllable devices in distribution networks [8]. The flexible dispatch region of a distribution network is summarized in a time-coupled power-energy band, taking into account the network topology  [9] and operational constraints. However, the proposed ITD framework does not consider the coordination between multiple TSOs in a data-preserving manner, and the proposed inner approximation is computationally inefficient, requiring solving multiple mixed-integer linear programming (milp)problems iteratively.
To enable privacy preservation and improve computational efficiency, distributed operation frameworks enable TSOs and DSOs to operate independently and collaborate effectively by sharing limited information with a subset of other operators [10, 11, 12, 13]. These proposed distributed frameworks can maintain data privacy and decision-making independence and are based on distributed AC OPF [11, 12] and MPOPF with receding horizon [13]. In addition to the aforementioned distributed algorithms, aladin is proposed for generic nonconvex optimization problems with convergence guarantees in [14]. aladin-type algorithms have been successfully applied to solve the single period AC OPF for heterogeneous power systems [15]. However, these aforementioned studies either lack a convergence guarantee or their scalability is limited by the computational complexity, which so far hinders an application to MPOPF in ITD systems.
@@
    In contrast to our previous work [3], we develop an aladin-type distributed nmpc approach for the multiperiod coordination of ITD systems within the proposed real-time framework. This approach is capable of decoupling the large-scale dispatch problem associated with different system operators and individual time periods. This particular design distributes the computational complexity of MPOPF over different stakeholders for computational affordability while maintaining the privacy of relevant information. Furthermore, the measurement data (i.e., load profiles and solar and wind outputs) from the ENTSO-E Transparency Platform [17] is employed in a simulation, showcasing the proposed methodology’s effectiveness with over 100,000 state variables divided into 400 subproblems, demonstrating its scalability and practical applicability in TSO-DSO coordination.
@@ II-A1 Exact Feasible Set
where $e_{1}=[1,0,\cdots,0]^{\top}\in\mathbb{R}^{N_{\textrm{bus}}}$, $R=\textrm{{diag}}(\textrm{r})$, $X=\textrm{{diag}}(\textrm{x})$. Note that the $\alpha$-th element in vector $\mathbf{r}$ (resp. $\mathbf{x}$) denotes the resistance (resp.) reactance of branch $\alpha$, $U_{k}$ denotes the vector of squared voltage magnitude at the time instant $k$, $p^{\textrm{pcc}}_{k}$, $q^{\textrm{pcc}}_{k}$ denote active and reactive power exchanges with the transmission system at the PCC of the distribution system. We use vectors $P^{d}_{k}$, $Q^{d}_{k}$ to denote the active and reactive nodal power consumptions, $P^{l}_{k}$, $Q^{l}_{k}$ to denote the active and reactive branch flows, and $P^{s}_{k}$ to denote the nodal consumptions by distributed energy storage systems (ESS)at time period $k$. Moreover, (1a) fixes the voltage magnitude at the slack bus. Equations (1b)-(1d) are the LinDistFlow constraints. Upper and lower bounds (1f) restrict the voltage magnitude at each bus and the charging/discharging power of ESS s. We rewrite the above power flow equations (1a)-(1d) in a compact form:
@@
The proof can be found in the Appendix. As a result of the generalized proposition 1, for a given distribution grid, all the dependent variables in $\chi_{k}$ can be expressed explicitly in terms of the controllable power injections $P^{s}_{k}$, and thus, the exact feasible set is convex and can be written in a convex polytope with respect to $P^{s}_{k}$,
@@ II-A2 Maximum Volume Inner Approximation
Note that $\overline{P}^{\textrm{appr}}_{k}$ and $\underline{P}^{\textrm{appr}}_{k}$ are upper and lower bounds of the inner hyperbox approximation. To maximize the performance of the resulting ESS system, we adopt the so-called maximum volume inner hyperbox [22]. The hyperbox (5) can be written as $\mathcal{B}^{s}_{k}(\zeta,\zeta+\xi)$ and the inner approximation can be obtained by solving the following optimization problem:
@@ II-B2 Multiperiod AC Optimal Power Flow
We utilize MPOPF for coordinating different TSOs at the transmission level, i.e.,
@@
$P^{b}_{k}(\Theta_{k},V_{k})$ $=C^{g}P^{g}_{k}-P^{d}_{k}-{\color[rgb]{0,0,0}C^{\textrm{pcc}}P^{\textrm{pcc}}_{k}}-C^{s}P^{s}_{k},$ (9b)
$Q^{b}_{k}(\Theta_{k},V_{k})$ $=C^{g}Q^{g}_{k}-Q^{d}_{k}-{\color[rgb]{0,0,0}C^{\textrm{pcc}}Q^{\textrm{pcc}}_{k}},$ (9c)
@@
where $P^{b}_{k}$, $Q^{b}_{k}$ represent the vectors of active and reactive power injections for all buses at time period $k$, expressed as nonlinear mappings $\mathbb{R}^{N_{\textrm{bus}}}\times\mathbb{R}^{N_{\textrm{bus}}}\mapsto\mathbb{R}^{N_{\textrm{bus}}}$ with respect to nodal voltage angles $\Theta_{k}$ and magnitudes $V_{k}$. Similarly, $S^{l}_{k}$ are nonlinear mappings $\mathbb{R}^{N_{\textrm{bus}}}\times\mathbb{R}^{N_{\textrm{bus}}}\mapsto\mathbb{R}^{N_{\textrm{line}}}$ representing apparent branch power flows for all branches at time period $k$; for the detailed formulation of power injections and branch power flows, we refer readers to [23]. $P^{\textrm{pcc}}_{k}$, $Q^{\textrm{pcc}}_{k}$ stack active and reactive power exchanges (II-B1) for all distribution systems, respectively. Evidently, MPOPF (II-B2) constructs a simultaneous formulation of $N_{k}$ AC OPF problems with standard power flow constraints (9b)-(9e), coupled by intertemporal interactions (9f) (9g) and the corresponding upper and lower bounds (9h). Notice that ESS s possess time-variant power limits in (9h), due to the inner hyperbox approximation (5) utilized for aggregating distribution systems.
@@ III Methodology
This section presents the proposed distributed real-time coordination framework of ITD systems using a receding horizon scheme while considering day-ahead forecast and actual values. Compared to the classical distributed MPC scheme, only solving the structured optimal control problem either in a spatially distributed manner or in a temporally distributed manner, our approach decouples the optimization problems across both different system operators and periods, with each subproblem representing an individual single-period AC OPF of a single transmission system.
@@ III-A Distributed Formulation
$\sum_{k\in\mathcal{K}}\Lambda_{k,\ell}x_{k,\ell}={\color[rgb]{0,0,0}0},\;\ell\in\mathcal{R},$
where the sparse matrices $\Lambda_{k,\ell}$ contain non-zero elements of $\{-1,1,\Delta t\}$, connecting the state variables current $E_{k,\ell}$ and $P^{g}_{k,\ell}$ with neighboring time periods $\{k-1,k+1\}$ for each transmission system $\ell$. Regarding spatial coupling among different TSOs, we follow the idea of sharing components [24], i.e., sharing nodal voltage angles and magnitudes at both sides of connecting tie-lines between neighboring transmission systems. The resulting spatial coupling constraints are linear and can be written in the following affinely coupled form
$\sum_{\ell\in\mathcal{R}}\Gamma_{k,\ell}x_{k,\ell}={\color[rgb]{0,0,0}0},\;k\in\mathcal{K},$
@@ III-B Real-Time Distributed Coordination Scheme
Based on the local measurements collected in Step 1), Step 2) of Algorithm 1 outlines the tailored ALADIN algorithm to solve (III-A). Step 2.a) solves $N_{k}\cdot|\mathcal{R}|$ subproblems, in which the regional TSO deals with the $N_{k}$ temporal subproblems in parallel locally. These problems are constructed using the Lagrangian of (III-A) by dualizing the spatial coupling (10f) and temporal coupling (10i). Based on the decoupled solutions $y_{k,\ell}$, Step 2.b) computes sensitivities of objective and constraints with respect to the current iteration of ALADIN. Here, in order to improve the numerical robustness of the algorithm, a small perturbation is added to the second-order derivatives (13) approximated by positive definite $H_{k,\ell}$ Notice that under a mild assumption for the perturbation as outlined in [3, Theorem 2], the local quadratic convergence can be guaranteed. Step 2.c) solves the coupled QP (2c) with only equality constraints. Taking the temporal coupling (14c) as local equality constraints for the $\ell$-th region, one can solve (2c) in a decentralized manner that only requires neighbor-to-neighbor communications. For more details, the reader is referred to [25]. A practical stopping criterion is to stop if the following conditions hold,
$\max_{k\in\mathcal{K}}\left\|\sum_{\ell\in\mathcal{R}}\Gamma_{k,\ell}y_{k,\ell}\right\|\leq\epsilon,\;\max_{\ell\in\mathcal{R}}\left\|\sum_{k\in\mathcal{K}}\Lambda_{k,\ell}y_{k,\ell}\right\|\leq\epsilon.$
After local solutions are determined at the transmission level, the inputs $(P_{0,\ell}^{s,*},\Delta P_{0,\ell}^{g,*})$ are deployed to the respective generators and storages. Step 4) in Algorithm 1 is a warm-start for Step 2) for the next online iteration following the strategy proposed in [26].

### 20 Feb 2024 v2

@@ Abstract
This paper proposes a real-time distributed operational architecture to coordinate integrated transmission and distribution systems (ITD). At the distribution system level, the distribution system operator (DSO)calculates the aggregated flexibility of all controllable devices by power-energy envelopes and provides them to the transmission system operator (TSO). At the transmission system level, a distributed nonlinear model predictive control (nmpc)approach is proposed to coordinate the economic dispatch of multiple TSOs, considering the aggregated flexibility of all distribution systems. The subproblems of the proposed approach are associated with different TSOs and individual time periods. In addition, the aggregated flexibility of controllable devices in distribution networks is encapsulated, re-calculated, and communicated through the power-energy envelopes, facilitating a reduction in computational complexity and eliminating redundant information exchanges between TSOs and DSOs, thereby enhancing privacy and security. The framework’s effectiveness and applicability in real-world scenarios are validated through simulated operational scenarios on a summer day in Germany, highlighting its robustness in the face of significant prediction mismatches due to severe weather conditions.
@@
Submitted to the 23nd Power Systems Computation Conference (PSCC 2024).
This work was supported in part by the BMBF-project ENSURE II with grant number 03SFK1F0-2, in part by the Swiss National Science Foundation (SNSF) under the NCCR Automation project, grant agreement 51NF40_180545, and in part by the Swiss Federal Office of Energy’s “SWEET” programme and performed in the PATHFNDR consortium. (Corresponding author: Yuning Jiang)
@@ I Introduction
Multiperiod dispatch problems for ITD systems usually couple individual steady-state optimal power flow (OPF)optimization problems over multiple time periods [4, 5, 6]. The coupling constraints include the generator ramping limits, the model of distributed energy storage systems (ESS), and other time-dependent constraints to consider the controllable devices with time-variant properties. However, it is still a challenge to solve a multiperiod AC optimal power flow (MPOPF)for ITD systems. On the one hand, the AC OPF is generally NP-hard [7], and the complexity of solving an MPOPF is further magnified by the intercoupling of subsequent time periods [5]. On the other hand, collecting necessary and realistic data from multiple stakeholders (i.e., TSOs and DSOs), including grid topology, load profiles, and other sensitive information regarding consumer behaviors, is either not preferred or restricted by regulations [3]. To address these challenges and achieve an efficient operation of the overall ITD, recent research analyzed the determination of the aggregated flexibility of the controllable devices in distribution networks [8, 9]. The flexible dispatch region of a distribution network is summarized in a time-coupled power-energy band, taking into account the network topology [10] and operational constraints. However, the proposed ITD framework does not consider the coordination between multiple TSOs in a data-preserving manner, and the proposed inner approximation is computationally inefficient, requiring solving multiple mixed-integer linear programming (milp)problems iteratively.
To enable privacy preservation and improve computational efficiency, distributed operation frameworks enable TSOs and DSOs to operate independently and collaborate effectively by sharing limited information with a subset of other operators [11, 12, 13, 14, 15, 16]. These proposed distributed frameworks can maintain data privacy and decision-making independence and are based on distributed AC OPF [17, 12] and MPOPF with receding horizon [18]. In addition to the aforementioned distributed algorithms, aladin is proposed for generic nonconvex optimization problems with convergence guarantees in [19]. aladin-type algorithms have been successfully applied to solve the single period AC OPF for heterogeneous power systems by a single-machine numerical simulation [20, 21, 15], as well as in a geographically distributed environment [22]. However, these aforementioned studies either lack a convergence guarantee or their scalability is limited by the computational complexity, which so far hinders an application to MPOPF in ITD systems.
@@
    In contrast to our previous work [3], we develop an aladin-type distributed nmpc approach for the multiperiod coordination of ITD systems within the proposed real-time framework. This approach is capable of decoupling the large-scale dispatch problem associated with different system operators and individual time periods. This particular design distributes the computational complexity of MPOPF over different stakeholders for computational affordability while maintaining the privacy of relevant information.
  3. 3.
    We conduct a comprehensive simulation by using real-world measurement data—including load profiles and solar and wind outputs—from a summer day in Germany marked by significant prediction discrepancies due to heavy rainfall, sourced from the ENTSO-E Transparency Platform [^1] The data utilized in this paper is available online at the ENTSO-E Transparency Platform: https://transparency.entsoe.eu [24]. This simulation, involving over 100,000 state variables divided into 400 subproblems at the transmission level, underscores the proposed approach’s efficiency, scalability, and practical relevance for TSO-DSO coordination.
@@
The rest of this paper is organized as follows: Section II presents the system model and problem formulation. Section III introduces the proposed distributed algorithm with implementation details. Section IV elaborates on numerical results. Section V concludes this paper.
@@ II-A1 Exact Feasible Set
where $e_{1}=[1,0,\cdots,0]^{\top}\in\mathbb{R}^{N_{\textrm{bus}}}$, $R=\textrm{{diag}}(\textrm{r})$, $X=\textrm{{diag}}(\textrm{x})$. $\mathbf{r},\,\mathbf{x}\in\mathbb{R}^{N_{\text{line}}}$ denote the resistance and reactance vectors respectively. $U_{k}$ denotes the vector of squared voltage magnitude at the time instant $k$, $p^{\textrm{pcc}}_{k}$, $q^{\textrm{pcc}}_{k}$ denote active and reactive power exchanges with the transmission system at the PCC of the distribution system. We use vectors $P^{d}_{k}$, $Q^{d}_{k}$ to denote the active and reactive nodal power consumptions, $P^{l}_{k}$, $Q^{l}_{k}$ to denote the active and reactive branch flows, and $P^{s}_{k}$ to denote the nodal consumptions by distributed energy storage systems (ESS)at time period $k$. Moreover, (1a) fixes the voltage magnitude at the slack bus. Equations (1b)-(1d) are the LinDistFlow constraints. Upper and lower bounds (1f) restrict the voltage magnitude at each bus and the charging/discharging power of ESS s. We rewrite the above power flow equations (1a)-(1d) in a compact form:
@@
The detailed proof can be found in the Appendix. As a result of the generalized proposition 1, for a given distribution grid, all the dependent variables in $\chi_{k}$ can be expressed explicitly in terms of the controllable power injections $P^{s}_{k}$, and thus, the exact feasible set is convex and can be written in a convex polytope with respect to $P^{s}_{k}$,
@@ II-A2 Maximum Volume Inner Approximation
Note that $\overline{P}^{\textrm{appr}}_{k}$ and $\underline{P}^{\textrm{appr}}_{k}$ are upper and lower bounds of the inner hyperbox approximation. The 2-dimensional green box in Fig. 2 represents an inner hyperbox approximation to the exact feasible set (blue polytope). To maximize the performance of the resulting ESS system, we adopt the so-called maximum volume inner hyperbox [29]. The hyperbox (5) can be written as $\mathcal{B}^{s}_{k}(\zeta,\zeta+\xi)$ and the inner approximation can be obtained by solving the following optimization problem:
@@ II-B2 Multiperiod AC Optimal Power Flow
The bus injection model [30] with complex voltages expressed in polar coordinates is employed at the transmission level. Here, $\Theta_{k},\,V_{k}$ stack nodal voltage angles $\theta_{k,i}$ and magnitudes $v_{k,i}$ for all bus $i$ at time period $k$ respectively. $P^{\textrm{pcc}}_{k}$, $Q^{\textrm{pcc}}_{k}$ stack active and reactive power exchanges (II-B1) for all distribution systems, respectively. $Y=G+\mathbf{j}B$ denote the complex bus admittance matrix, where $\mathbf{j}=\sqrt{-1}$ and $G,B\in\mathbb{R}^{N_{\text{bus}}\times N_{\text{bus}}}$.
The resulting multiperiod AC optimal power flow (MPOPF)for coordinating different TSOs can be written as
@@
$P^{b}_{k}(\Theta_{k},V_{k})$ $=C^{g}P^{g}_{k}-P^{d}_{k}-C^{\textrm{pcc}}P^{\textrm{pcc}}_{k}-C^{s}P^{s}_{k},$ (9b)
$Q^{b}_{k}(\Theta_{k},V_{k})$ $=C^{g}Q^{g}_{k}-Q^{d}_{k}-C^{\textrm{pcc}}Q^{\textrm{pcc}}_{k},$ (9c)
@@
where $P^{b}_{k},Q^{b}_{k}:\,\mathbb{R}^{N_{\textrm{bus}}}\times\mathbb{R}^{N_{\textrm{bus}}}\mapsto\mathbb{R}^{N_{\textrm{bus}}}$ represent the vector functions of active and reactive power injections for all buses at time period $k$, and the corresponding $i$-th element can be expressed as
$\left[P^{b}_{k}\right]_{i}=v_{k,i}\sum_{j\in\mathcal{N}}v_{k,j}\left(G_{ij}\cos\theta_{k,ij}+B_{ij}\sin\theta_{k,ik}\right),$
$\left[Q^{b}_{k}\right]_{i}=v_{k,i}\sum_{j\in\mathcal{N}}v_{k,j}\left(G_{ij}\sin\theta_{k,ij}-B_{ij}\cos\theta_{k,ij}\right),$
with angle difference $\theta_{k,ij}=\theta_{k,i}-\theta_{k,j}$. Similarly, $S^{l}_{k}$ are nonlinear mappings $\mathbb{R}^{N_{\textrm{bus}}}\times\mathbb{R}^{N_{\textrm{bus}}}\mapsto\mathbb{C}^{N_{\textrm{line}}}$ representing apparent branch power flows for all branches at time period $k$; for the detailed formulation of branch power flows, we refer readers to [30]. Evidently, MPOPF (II-B2) constructs a simultaneous formulation of $N_{k}$ AC OPF problems with standard power flow constraints (9b)-(9e), coupled by intertemporal interactions (9f) (9g) and the corresponding upper and lower bounds (9h). Notice that ESS s possess time-variant power limits in (9h), due to the inner hyperbox approximation (5) utilized for aggregating distribution systems.
@@ III Methodology
This section presents the proposed distributed real-time coordination framework of ITD systems using a receding horizon scheme while considering day-ahead forecast and actual values. Compared to the classical distributed MPC scheme, only solving the structured optimal control problem either in a spatially distributed manner or in a temporally distributed manner [31], our approach decouples the optimization problems across both different system operators and periods, with each subproblem representing an individual single-period AC OPF of a single transmission system.
@@ III-A Distributed Formulation
$\sum_{k\in\mathcal{K}}\Lambda_{k,\ell}x_{k,\ell}=0,\;\ell\in\mathcal{R},$
where the sparse matrices $\Lambda_{k,\ell}$ contain non-zero elements of $\{-1,1,\Delta t\}$, connecting the state variables current $E_{k,\ell}$ and $P^{g}_{k,\ell}$ with neighboring time periods $\{k-1,k+1\}$ for each transmission system $\ell$.
Regarding spatial coupling among different TSOs, we follow the idea of sharing components [32], i.e., sharing nodal voltage angles and magnitudes at both sides of connecting tie-lines between neighboring transmission systems. The resulting spatial coupling constraints are linear and can be written in the following affinely coupled form
$\sum_{\ell\in\mathcal{R}}\Gamma_{k,\ell}x_{k,\ell}=0,\;k\in\mathcal{K},$
@@ III-B Real-Time Distributed Coordination Scheme
Based on the local measurements collected in Step 1), Step 2) of Algorithm 1 outlines the tailored ALADIN algorithm to solve (III-A). Step 2.a) solves $N_{k}\cdot|\mathcal{R}|$ subproblems, in which the regional TSO deals with the $N_{k}$ temporal subproblems in parallel locally. These problems are constructed using the Lagrangian of (III-A) by dualizing the spatial coupling (10f) and temporal coupling (10i). Based on the decoupled solutions $y_{k,\ell}$, Step 2.b) computes sensitivities of objective and constraints with respect to the current iteration of aladin. Here, in order to improve the numerical robustness of the algorithm, a small perturbation is added to the second-order derivatives (13) approximated by positive definite $H_{k,\ell}$ Notice that under a mild assumption for the perturbation as outlined in [3, Theorem 2], the local quadratic convergence can be guaranteed. Step 2.c) solves the coupled QP (2c) with only equality constraints. Taking the temporal coupling (14c) as local equality constraints for the $\ell$-th region, one can solve (2c) in a decentralized manner that only requires neighbor-to-neighbor communications. For more details, the reader is referred to [33]. Algorithm 1 terminates if the primal conditions
$\max_{k\in\mathcal{K}}\left\|\sum_{\ell\in\mathcal{R}}\Gamma_{k,\ell}y_{k,\ell}\right\|\leq\epsilon,\,\max_{\ell\in\mathcal{R}}\left\|\sum_{k\in\mathcal{K}}\Lambda_{k,\ell}y_{k,\ell}\right\|\leq\epsilon,$ (15a)
and dual condition
$\max_{\begin{subarray}{c}k\in\mathcal{K}\\ \ell\in\mathcal{R}\end{subarray}}\left\lVert y_{k,\ell}-x_{k,\ell}\right\rVert\leq\epsilon$ (15d)
hold. Practically, the dual condition (15d) is sufficient to ensure a small violation of the condition (III-B), when the predefined tolerance $\epsilon$ is small enough [34]. Under some regularity assumptions, Algorithm 1 has local quadratic convergence guarantees for both primal and dual iterations. One can construct the proof of this result by following that the coupled QP (2c) is equivalent to the Newton-type method while the local solutions maps are Lipschitz continuous. A detailed analysis can be found in [19, 20].
When employing Algorithm 1 as an online solver, Algorithm 1 presents a receding horizon scheme to coordinate the ITD system in the closed loop. During the online process, each local TSO measures the states $(E_{0,\ell},P_{0,\ell}^{g})$ and then, Algorithm 1 solves (III-A) in a distributed manner. After local solutions are determined at the transmission level, the determined inputs $(P_{1,\ell}^{s,*},\Delta P_{1,\ell}^{g,*})$ are allocated to the respective generators and storages. Notably, step 4 in Algorithm 1 serves as an initialization phase for step 2 in the ensuing online cycle, adhering to the methodology outlined in [35].